# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

### v3

@@ I Introduction
However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [41, 39], as illustrated in Figure 1. This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted (represented by the black lines), the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext. This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference. Notably, this risk is amplified in the emerging paradigm of confidential Model-as-a-Service (MaaS). Here, TEE adoption for inference privacy [5, 10] introduces a critical architectural trade-off: high-throughput architectures require deliberately externalizing the massive KV-cache from the TEE’s protection boundary. This architectural design—not a vulnerability—directly exposes the KV-cache to the Cloud Service Provider (CSP), who inherently possesses it without requiring any additional exploit [43].
Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with the architectural realities that expose it in plaintext, creates a severe and practical attack vector. A critical research gap therefore exists in mitigating these risks, necessitating a practical and secure defense mechanism.
@@
We first explore two relatively direct attack paths. The first is the KV-cache Inversion Attack, a method that attempts to leverage known model weight matrices to directly reverse-calculate the input from the KV-cache. The second, more general-purpose approach is the KV-cache Collision Attack. This attack reframes input reconstruction as a matching task based on forward computation: an adversary iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match. Because this method does not rely on any reverse computation, it has broader applicability. Finally, we introduce a novel, semantic-based KV-cache Injection Attack. The core of this attack is to leverage the LLM’s powerful capability to understand and execute instructions. By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an adversary can induce the model to “echo” or generalize the core information held within the KV-cache.
@@
Therefore, we propose KV-Cloak, a lightweight and secure mechanism for KV-cache obfuscation designed to overcome these challenges. At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy. Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block. This dynamic permutation prevents adversaries from building stable algebraic relations across multiple queries. To further enhance performance, KV-Cloak utilizes operator fusion, an optimization where a portion of the secret obfuscation matrices is algebraically fused into the LLM’s attention layer weights offline. This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
@@ II-A Transformer-based LLM Inference
KV-cache in LLM Inference. To mitigate the computational redundancy of autoregressive generation, LLMs utilize the KV-cache mechanism [28]. Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states in memory. When generating a new token $x_{i}$, the model only computes the current $q_{i},k_{i},v_{i}$, performs attention between $q_{i}$ and the cached history, and subsequently appends the new $(k_{i},v_{i})$ pair to the cache. This strategy significantly reduces latency but necessitates managing a persistent state proportional to the sequence length.
@@ III-B3 KV-cache Injection Attack
Strategic Advantages: Robustness and Efficiency. This attack offers distinct advantages in robustness compared to the Collision Attack. Specifically, it remains effective against KV-cache eviction strategies like H2O [46]. While such lossy compression breaks the strict mathematical correspondence required for algebraic attacks, it typically preserves the semantic gist of the context. The Injection Attack successfully exploits these residual semantic vectors to hallucinate or reconstruct private data. Furthermore, the attack is highly efficient, requiring only a single generation pass. This vector underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
@@ VII Discussion and Future Work
Key Management and Hardware Security Integration. The security of KV-Cloak fundamentally relies on the confidentiality of its (megabyte-scale) key matrices. Currently, we assume these secrets are protected via TEEs. Future work should explore the deep integration of KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture. Furthermore, to counter long-term cryptanalysis in persistent services, developing efficient, low-latency key rotation protocols is essential.
@@
Adaptation to Quantized Models. Our current prototype targets floating-point models. As the industry pivots toward integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak is a priority. This necessitates designing new, lossless reversible transformations suitable for discrete data types, potentially based on mathematical structures like modular arithmetic.

### v4

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9919276-8-0 https://dx.doi.org/10.14722/ndss.2026.240258 www.ndss-symposium.org
@@ I Introduction
However, the storage and potential sharing of the KV-cache introduce significant yet underexplored privacy concerns [41, 39], as illustrated in Figure 1. This vulnerability stems from a critical trade-off made in production systems: while end-user communication with the LLM service is typically encrypted, the KV-cache itself is almost always processed, transmitted between compute nodes, and persisted in plaintext. This design choice is a concession to performance, as the unacceptable latency overhead from cryptographically securing the often gigabyte-scale KV-cache would violate the stringent demands of real-time inference. Notably, this risk is amplified in the emerging paradigm of confidential Model-as-a-Service (MaaS). Here, TEE adoption for inference privacy [5, 10] introduces a critical architectural trade-off: high-throughput architectures require deliberately externalizing the massive KV-cache from the TEE’s protection boundary. This architectural design—not a vulnerability—directly exposes the KV-cache to the Cloud Service Provider (CSP), who inherently possesses it without requiring any additional exploit [43].
Crucially, the direct correlation between the KV-cache and user inputs [7, 21], combined with its architectural exposure in plaintext, creates a severe and practical attack vector. A critical research gap therefore exists in mitigating these risks, necessitating a practical and secure defense mechanism.
@@
We first explore two relatively direct attack paths. The first is the KV-cache Inversion Attack, leveraging known model weight matrices to directly reverse-calculate the input from the KV-cache. The second, more general-purpose approach is the KV-cache Collision Attack. This attack reframes input reconstruction as a matching task based on forward computation: an adversary iteratively uses a local model instance to generate KV-caches for candidate inputs and compares them against the intercepted target KV-cache to find a match. Because this method does not rely on any reverse computation, it has broader applicability. Finally, we introduce a novel, semantic-based KV-cache Injection Attack. This attack leverages the LLM’s powerful capability to understand and execute instructions. By appending a specific instruction, such as “Repeat the previous content,” to the end of an intercepted KV-cache context, an adversary can induce the model to “echo” or generalize the core information held within the KV-cache.
@@
Therefore, we propose KV-Cloak, a lightweight and secure KV-cache obfuscation mechanism addressing these challenges. At its core, KV-Cloak employs a reversible matrix-based obfuscation scheme that guarantees lossless model accuracy. Its security is multi-layered: it applies secret invertible linear transformations to obscure statistical properties, and critically, introduces a one-time random permutation matrix for each data block. This dynamic permutation prevents adversaries from building stable algebraic relations across multiple queries. To further enhance performance, KV-Cloak utilizes operator fusion, algebraically fusing a portion of the secret obfuscation matrices into the LLM’s attention layer weights offline. This shifts the primary computational cost away from the latency-sensitive online inference phase, striking an effective balance between robust security, lossless accuracy, and high performance.
@@ II-A Transformer-based LLM Inference
KV-cache in LLM Inference. To mitigate the computational redundancy of autoregressive generation, LLMs utilize the KV-cache mechanism [28]. Instead of recalculating the Key ($K$) and Value ($V$) matrices for the entire context window at each step, the system persists these intermediate states. When generating a new token $x_{i}$, the model only computes the current $q_{i},k_{i},v_{i}$, performs attention between $q_{i}$ and the cached history, and subsequently appends the new $(k_{i},v_{i})$ pair to the cache. This strategy significantly reduces latency but necessitates managing a persistent state proportional to the sequence length.
@@ III-B3 KV-cache Injection Attack
Strategic Advantages: Robustness and Efficiency. This attack offers distinct advantages in robustness compared to the Collision Attack. Specifically, it remains effective against KV-cache eviction strategies like H2O [46]. While such lossy compression breaks the strict mathematical correspondence required for algebraic attacks, it typically preserves the semantic gist of the context. The Injection Attack successfully exploits these residual semantic vectors to hallucinate or reconstruct private data. Furthermore, the attack is highly efficient, requiring only a single generation pass. This attack underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
@@ VII Discussion and Future Work
Key Management and Hardware Security Integration. The security of KV-Cloak fundamentally relies on the confidentiality of its (megabyte-scale) key matrices. Currently, we assume these secrets are protected via TEEs. Future work should deeply integrate KV-Cloak with hardware-level security mechanisms, such as TEEs or confidential GPUs, to construct a more robust defense-in-depth architecture. Furthermore, to counter long-term cryptanalysis in persistent services, developing efficient, low-latency key rotation protocols is essential.
@@
Adaptation to Quantized Models. Our current prototype targets floating-point models. As the industry shifts to integer quantization (e.g., INT8 or INT4) for efficiency, extending KV-Cloak becomes a priority. This requires designing new lossless reversible transformations for discrete data types, potentially based on structures like modular arithmetic.
@@ VIII Conclusion
Acknowledgment
This research is supported in part by the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01169), and the National Natural Science Foundation of China under Grants (62441238, U2441240).